Two pellets, each with a charge of 1.2 microcoulombs (1.2×10−6 C ), are located 3.6 cm (3.6×10−2 m ) apart. Find the electric force between them.
Answer #1

here,

charge on 1 , q1 = 1.2 uC = 1.2 * 10^-6 C

charge on 2 , q2 = q1

distance , r = 2.6 cm = 0.036 m

the electric force between them , F = k * q1 * q2 /r^2

F = 9 * 10^9 * 1.2 * 10^-6 * 1.2* 10^-6 /0.036^2 N

F = 10 N

the electric force between them is 10 N